# AN OPEN SUBSET OF THE VARIETY OF $n$- DIMENSIONAL ALGEBRAS, CLASSIFICATION AND AUTOMORPHISMS Ural Bekbaev

### v2

@@ Abstract
Classification, up to isomorphism, of algebras from a non-empty subset of the variety of $n$- dimensional algebras is presented. It is shown that these algebras have only trivial automorphism and if the basic field is algebraically closed then it is an open dense subset of the variety of $n$- dimensional algebras.
@@ 1. Introduction
In this paper we provide another algorithm for the construction of that matrix $P(A)$, show that it works well in arbitrary $n$ and $\mathbb{F}$ case, the set of algebras, for which $P(A)$ is invertible ($\mathcal{A}_{00}(n)$), is not empty if even $\mathbb{F}$ is not algebraically closed. We classify all algebras for which $P(A)$ is invertible, provide a formulae to find a canonical form of any such $A$, show that the automorphism of $A$ is trivial and it does not have invertible derivation. Unlike [1] we show that, if $\mathbb{A}$ is $n$-dimensional, $\mathbb{A^{\prime}}$ is $n^{\prime}$-dimensional algebras then $\mathbb{A}\oplus\mathbb{A^{\prime}}\in\mathcal{A}_{00}(n+n^{\prime})$ if and only if $\mathbb{A}\in\mathcal{A}_{00}(n)$ and $\mathbb{A^{\prime}}\in\mathcal{A}_{00}(n^{\prime})$.
@@ 3. The main result
If $A\in\mathcal{A}(n)$, $A^{\prime}\in\mathcal{A}(n^{\prime})$ then $\mathbb{D}=\mathbb{A}\oplus\mathbb{A^{\prime}}\in\mathcal{A}(n+n^{\prime})$ if and only if $A\in\mathcal{A}_{00}(n)$, $A^{\prime}\in\mathcal{A}_{00}(n^{\prime})$.
@@
Let $A=(A_{1}|A_{2})=\begin{pmatrix}0&1&1&0\\ 1&0&0&-1\end{pmatrix}$. It is a simple algebra, but it is not in $\mathcal{A}_{00}(2)$, $P(A)$ is singular as far as $\overline{tr}(A)=0$.

### v3

@@ Abstract
Classification, up to isomorphism, of algebras from a non-empty subset of the variety of $n$- dimensional algebras is presented. It is shown that these algebras have only trivial automorphism and if the basic field is algebraically closed then it is an open dense subset of the variety of $n$- dimensional algebras. Some nice properties of these subsets are considered as well.
@@ 1. Introduction
In this paper we provide another algorithm for the construction of that matrix $P(A)$, show that it works well in arbitrary $n$ and $\mathbb{F}$ case, the set of algebras, for which $P(A)$ is invertible ($\mathcal{A}_{00}(n)$), is not empty if even $\mathbb{F}$ is not algebraically closed. We classify all algebras for which $P(A)$ is invertible, provide a formulae to find a canonical form of any such $A$, show that the automorphism group of $A$ is trivial and it does not have invertible derivation. Unlike [1] we show that, if $\mathbb{A}$ is $n$-dimensional, $\mathbb{A^{\prime}}$ is $n^{\prime}$-dimensional algebras then $\mathbb{A}\oplus\mathbb{A^{\prime}}\in\mathcal{A}_{00}(n+n^{\prime})$ if and only if $\mathbb{A}\in\mathcal{A}_{00}(n)$ and $\mathbb{A^{\prime}}\in\mathcal{A}_{00}(n^{\prime})$. Moreover, it is shown that $\mathbb{A}\otimes\mathbb{A^{\prime}}\in\mathcal{A}_{00}(nn^{\prime})$ if and only if $\mathbb{A}\in\mathcal{A}_{00}(n)$ and $\mathbb{A^{\prime}}\in\mathcal{A}_{00}(n^{\prime})$. A description of elements of $\mathbb{A}\in\mathcal{A}_{00}(2)$, up to isomorphism, is given as well.
@@ 3. The main result
If $\mathbb{A}\in\mathcal{A}(n)$, $\mathbb{A}^{\prime}\in\mathcal{A}(n^{\prime})$ then $\mathbb{D}=\mathbb{A}\oplus\mathbb{A^{\prime}}\in\mathcal{A}(n+n^{\prime})$ if and only if $A\in\mathcal{A}_{00}(n)$, $A^{\prime}\in\mathcal{A}_{00}(n^{\prime})$.
@@
Let $A=(A_{1}|A_{2})=\begin{pmatrix}0&1&1&0\\ 1&0&0&-1\end{pmatrix}$-a Jordan algebra. It is a simple algebra, but it is not in $\mathcal{A}_{00}(2)$, $P(A)$ is singular as far as $\overline{tr}(A)=0$.
Theorem 3.11.
If $\mathbb{A}\in\mathcal{A}(n)$, $\mathbb{A}^{\prime}\in\mathcal{A}(n^{\prime})$ then $\mathbb{D}=\mathbb{A}\otimes\mathbb{A^{\prime}}\in\mathcal{A}(nn^{\prime})$ if and only if $A\in\mathcal{A}_{00}(n)$, $A^{\prime}\in\mathcal{A}_{00}(n^{\prime})$.
Proof.
Indeed, if $e=(e_{1},e_{2},...,e_{n})$, $e^{\prime}=(e^{\prime}_{1},e^{\prime}_{2},...,e^{\prime}_{n^{\prime}})$ are corresponding bases then $e\otimes e^{\prime}$ is a basis for $\mathbb{D}$ and its MSC is
$D=A\otimes A^{\prime}=(A_{1}\otimes A^{\prime}_{1}|...|A_{1}\otimes A^{\prime}_{n^{\prime}}|A_{2}\otimes A^{\prime}_{1}|...|A_{2}\otimes A^{\prime}_{n^{\prime}}|...|A_{n}\otimes A^{\prime}_{1}|...|A_{n}\otimes A^{\prime}_{n^{\prime}}),$
$\overline{tr}(D)=\overline{tr}(A)\otimes\overline{tr}(A^{\prime})$, $tr((A_{i}\otimes A^{\prime}_{p})(A_{j}\otimes A^{\prime}_{q}))=tr(A_{i}A_{j}\otimes A^{\prime}_{p}A^{\prime}_{q})=tr(A_{i}A_{j})tr(A^{\prime}_{p}A^{\prime}_{q})$. It implies that $B(D)=B(A)\otimes B(A^{\prime})$ and $B(D)$ is non-singular if and only if $B(A),B(A^{\prime})$ are nonsingular. Similarly, $B(D^{o})=B(A^{o})\otimes B(A^{\prime o})$, $(B(A)\otimes B(A^{\prime}))^{-1}(B(A^{o})\otimes B(A^{\prime o}))=B(A)^{-1}B(A^{o})\otimes B(A^{\prime})^{-1}B(A^{\prime o})$, $P(D)=P(A)\otimes P(A^{\prime})$ and $P(D)$ is invertible if and only $P(A),P(A^{\prime})$ are invertible. ∎
Note that if $a\in\mathbb{F}$ is non-zero, $\mathbb{A}\in\mathcal{A}_{00}(n)$ then $a\mathbb{A}\in\mathcal{A}_{00}(n)$, where $a\mathbb{A}$ stands for the algebra with MSC $(aA_{1}|aA_{2}|...|aA_{n})$, as far as $P(aA)=aP(A)$. Therefore consideration of $\cup_{n=0}^{\infty}\mathcal{A}_{00}(n)$ with respect to $\oplus$,$\otimes$ , ⊗ and multiplication by a scalar operations is a natural thing, where $\mathcal{A}_{00}(0)$ stands for the zero element. Isn’t it an interesting object to explore?
One can use the classification theorem from [3] to describe elements of $\mathcal{A}_{00}(2)$ up to isomorphism. Here we present the corresponding result in $Char(\mathbb{F})\neq 2,3$ case and follow the notations of that paper.
Theorem 3.12.
Every element of $\mathcal{A}_{00}(2)$, in $Char(\mathbb{F})\neq 2,3$ case, is isomorphic to only one of the following algebras from $\mathcal{A}_{00}(2)$ given by theirs following MSC’s
  • $A_{1}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&1+\alpha_{2}&\alpha_{4}\\ \beta_{1}&-\alpha_{1}&1-\alpha_{1}&-\alpha_{2}\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{2},\alpha_{4},\beta_{1})\in\mathbb{F}^{4},4\alpha_{2}^{3}\beta_{1}+3\alpha_{1}^{2}\alpha_{2}^{2}-2\alpha_{1}\alpha_{2}^{2}+4\alpha_{2}^{2}\beta_{1}+\alpha_{2}^{2}+2\alpha_{1}^{2}\alpha_{2}-2\alpha_{1}\alpha_{2}+4\alpha_{1}^{2}\alpha_{4}-4\alpha_{1}^{3}\alpha_{4}-2\alpha_{1}\alpha_{4}-6\alpha_{1}\alpha_{2}\alpha_{4}\beta_{1}-2\alpha_{1}\alpha_{4}\beta_{1}-\alpha_{1}^{2}-\alpha_{4}^{2}\beta_{1}^{2}+2\alpha_{2}\alpha_{4}\beta_{1}\neq 0,$
    $\alpha_{1}-2\alpha_{1}^{2}+2\alpha_{1}^{3}+2\alpha_{1}\alpha_{2}-4\alpha_{1}^{2}\alpha_{2}+4\alpha_{1}^{3}\alpha_{2}+2\alpha_{1}\alpha_{2}^{2}-4\alpha_{1}^{2}\alpha_{2}^{2}+2\alpha_{1}\alpha_{4}-6\alpha_{1}^{2}\alpha_{4}+8\alpha_{1}^{3}\alpha_{4}+2\alpha_{1}\beta_{1}+6\alpha_{1}\alpha_{2}\beta_{1}+8\alpha_{1}\alpha_{2}^{2}\beta_{1}+4\alpha_{1}\alpha_{4}\beta_{1}-4\alpha_{1}^{2}\alpha_{4}\beta_{1}+4\alpha_{1}\alpha_{2}\alpha_{4}\beta_{1}\neq 0,$
  • $A_{2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&\beta_{2}&1-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3},\\ \alpha_{4}(2\alpha_{1}^{3}-2\alpha_{1}^{2}+2\alpha_{1}\beta_{2}^{2}+\alpha_{4}-2\beta_{2}^{2})(\alpha_{1}+\beta_{2}-1)\neq 0,$
  • $A_{4}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&\beta_{2}&1&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\beta_{1},\beta_{2})\in\mathbb{F}^{2},$
    $(\beta_{2}^{2}+2\beta_{1}+2\beta_{2}-1)(-2\beta_{2}^{2}+4\beta_{2}+5)\neq 0,$
  • $A_{6}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2},\\ \alpha_{4}(2(2\alpha_{1}^{2}-2\alpha_{1}+1)\alpha_{1}+\alpha_{4})\neq 0,$
  • $A_{8}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&1&0&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},$ $4\beta_{1}+1\neq 0$
Proof.
Proof consists of calculations of $B(A),P(A)$, finding out for which $A$ from the classification Theorem 1 [3] the inequalities $\det(B(A))\neq 0,\det(P(A))\neq 0$ are valid. ∎
@@
If $Char(\mathbb{F})\neq 2,3$ and $A\in\mathcal{A}_{00}(2)$ then $Der(A)=\{0\}$.
Example 3.16.
If $Char(\mathbb{F})\neq 2$ then for the algebra $A=\left(\begin{array}[]{cccc}1&0&1&-\frac{1}{2}\\ 0&-1&0&0\end{array}\right)$ one has $Aut(A)=\{I\}$, $Der(A)=\{0\}$, but $A$ is not in $\mathcal{A}_{00}(2)$.
In $Char(\mathbb{F})=2$ case the algebra $A=\left(\begin{array}[]{cccc}1&\alpha_{2}&1+\alpha_{2}&1\\ 1&-1&0&-\alpha_{2}\end{array}\right)$ has similar property: $Aut(A)=\{I\}$, $Der(A)=\{0\}$, but $A$ is not in $\mathcal{A}_{00}(2)$..
Corollary 3.17.